The small, repetitive hybrid events which have dominated the seismicity over the past week or so continued throughout the period. Amplitudes have been quite variable, and most are only being recorded at Gages, Long Ground and Windy Hill seismometers. Their rate of occurrence has averaged three to four per minute and maximum peak to peak amplitude on Gages has been in the range 10 to 15 mm throughout the period. A total of 14 larger events related to a separate process were recorded during this period - some correlate with big rock fall events from the lava dome, one of which was witnessed by scientists on the crater rim at about 17:00 on 26 January.
EDM measurements were carried out again today on the Long Ground - White's Yard - Castle Peak triangle today. The changes in line length continue to be within the error of the method. A new levelling line was marked out today in the Farrell's area; this will be occupied periodically and gives an additional method of measuring any deformation of the volcano.
Views of the new dome were again obtained today in partially clear conditions from the helicopter and from the crater rim at Chances Peak. Active growth continues on both the north and south sides of the new dome. Large blocks of lava are accumulating amongst smaller blocks and ash in the moat areas adjacent to the active areas of the dome. The northern area looks more stable than over the past few days and rock falls may start to decrease in number from this area over the next few days. Steam emissions continue from several parts of the crater.
Mr Wilkie Balgobin of the SRU arrived on island late on 26 January; he will be assisting with upkeep and rationalisation of the seismic and other electronic equipment used by MVO in their monitoring work.